# Enhancing LLMs for Governance with Human Oversight: Evaluating and Aligning LLMs on Expert Classification of Climate Misinformation for Detecting False or Misleading Claims about Climate Change

### v1

@@ 4.4 Expert annotation of claims in social media data
Two climate change communication experts, each with over 20 years of experience as professors and academic researchers on how climate science is communicated, annotated the sampled paragraphs from mostly engaging articles on social media (described in section 4.3) on the super-claim and sub-claim levels per the CARDS coding manualCoan et al. (2021).
@@ 5 Results
Fine-tuned GPT-3.5-turbo RoBERTalarge
Claim Code Precision Recall F1 Support Precision Recall F1 Support
0 0.94 0.94 0.94 491 0.76 0.96 0.85 491
1 0.71 1.00 0.83 24 0.71 0.62 0.67 24
2 0.93 0.93 0.93 14 1.00 0.79 0.88 14
3 0.57 1.00 0.73 8 0.57 0.50 0.53 8
4 0.95 0.94 0.95 274 0.93 0.63 0.75 274
5 0.99 0.89 0.94 103 0.90 0.62 0.74 103
Accuracy 0.94 914 0.81 914
Macro avg 0.85 0.95 0.88 914 0.81 0.69 0.74 914
Weighted avg 0.94 0.94 0.94 914 0.83 0.81 0.80 914
@@ 5.1 Evaluating the performance of LLMs on CARDS data
Using the zero-shot approach (described in section 4.1) to classify climate misinformation on CARDS test dataset, we encountered some inconsistencies in the way some LLMs are responding to our task. Figure 2 in the appendix shows the rate of valid responses per model. The prominent source of this inconsistency is the failure of models to comply with the instructions outlined in the prompt, resulting in responses that are irrelevant or with unstructured formatting. A similar inconsistent behavior of LLMs was also reported in prior research evaluating open-source LLMs on 20 NLP tasks Bavaresco et al. (2024). However, we observed instances of Llama3.1-8B failing to classify claims because it is "not sure if this is the right place to post this [request]" and that the model is "trying to get a better".
@@
Though GPT-4o is a much larger model compared to RoBERTalarge, in terms of parameter size, it is more budget friendly for stakeholders in climate change discourse such as researchers, policymakers, or think tanks to utilize this model when classifying climate misinformation as it does not require substantial investment in computational resources or expertise to develop a customized model such as CARDS for classifying climate misinformation. However, detecting false or misleading claims does not include the models’ capability to respond to or mitigate such claims unless they are augmented with relevant external knowledge as seen in use cases for fact-checking climate change claimsLeippold et al. (2024). Accordingly, future research is needed to explore and extend our work to determine whether models capable of detecting climate misinformation can augment general models, such as GPT-4o, with information about which claims are false or misleading to help orient these general models toward guard-railing against such claims.
@@
On the other hand, we observed a poor performance by the fine-tuned model in detecting claims related to climate impacts are not bad. Reviewing the annotated super-claims and sub-claims by experts, we found that the fine-tuned model is unable to accurately classify sub-claims within this category about the impacts of climate change on animal and plant species (see sub-claim 3.2 within the taxonomy of claims in Appendix A.2). We found that the fine-tuned GPT-3.5-turbo is biased toward inaccurately classifying claims regarding climate change impacts on animal and plant species as false, possibly due to biases in the CARDS training data that we originally fine-tuned the model on (see Limitations in section Limitations). This indicates additional fine-tuning is needed to enhance the ability of the model to differentiate between text describing positive versus negative impacts of climate change.
@@ 6 Conclusion
As developers and researchers test the potential of LLMs to persuade and misinform at scale Matz et al. (2024); Zhang et al. (2024), evaluating the potential for LLMs to be part of the solution for governing online dis/misinformation rather than the problem becomes a task of great importance. In this context, the overarching goal of this paper was to demonstrate the crucial role of human-oversight when leveraging LLMs for governance by evaluating the performance of, both open-source and proprietary, LLMs in the application of classifying false or misleading claims about climate change. We also aimed to compare (1) how these models performed against existing tools designed for the same task, and (2) with the assessments made by climate-communication and mis-information experts.
@@
Another limitation is the reliance of our research on a single taxonomy of claims developed by the authors of CARDS Coan et al. (2021) that was leveraged to fine-tune GPT-3.5.-turbo. This introduces two sets of biases. First, as the CARDS annotated dataset was based on an expert review of climate skeptic and contrarian domains, whether the GPT-3.5-turbo model is capable of precisely discriminating between accurate and inaccurate climate change claims within high credible sources (e.g. The New York Times, CNN, etc.) is an open question. One way to address this problem, as the authors of the CARDS model have recently moved toward, is a two-stage approach for classifying claims that first determines the veracity of the claim and in the second stage labels the category of false claims Rojas et al. (2024). A next step, therefore, is to benchmark GPT-3.5-turbo model’s performance in classifying claims from high credible sources for comparison to the updated CARDS model and continue fine-tuning as necessary on text sourced from domains with varying credibility.
A second model bias is the inability to classify claims in social media posts that do not fall within the CARDS taxonomy. For instance, the original taxonomy upon which CARDS is based includes claims about human health impacts as a category but was excluded in the final model due it its low prevalence Coan et al. (2021). However, this may be a function of the ideological-skew of climate skeptic blogs on which CARDS was trained that ignored this dimension of climate change impacts and/or temporal trends increasing the prominence of health impacts. As a result, climate change communication experts annotating the social media claims observed a substantial number of false claims about the health impacts of climate change on humans that did not fall within the current CARDS model and which the GPT 3.5-Turbo was unable to classify.
In addition, the model’s poor performance in classifying claims about the impacts of climate change on animal and plant species (see Table 1), could also be attributed to the under-representation of these examples in the CARDS dataset used for fine-tuning. Moving forward, fine-tuning the GPT-3.5-turbo model on additional expert annotated datasets, for example from Climate Feedback  [^10] https://science.feedback.org would likely enhance the model’s performance in accurately classifying a wider range of claims. These limitations stress the importance of benchmarking the performance of LLMs against data collected from "the wild" as we did in this paper and fine-tuning accordingly to ensure optimal performance in detecting misinformation online.
@@ A.1 Climate Change Keywords
A full list of the compiled climate change keywords identified by climate experts that are used to scrape and filter relevant climate change articles: climate change - climate crisis - climate effects - climate hoax - climate policy - climate resilience - climate science - climate summit - global warming - greenhouse gas - greenhouse gases - IPCC - green energy - climate hypocrisy - paris agreement - paris climate - net zero - net-zero - COP26 - climate conversation - climate test - climate gap - climate activists - climate activist - clean energy - climate negotiations - climate deal - green new deal - climate conference - green technology - green tech - climate fearmongering - climate fears - climate anxiety - carbon capture
@@ A.4 System Prompt
CARDS: Computer Assisted Recognition of Denial and Skepticism, is a machine learning project.
Our aim is to train a computer to automatically detect and categorize misinformation about climate
change. The end goal is that a computer can look at some text and successfully identify any climate
misinformation - and even identify specific denialist claims. If successful, this will enable us to
travel back in time and build a history of climate misinformation, including when myths originated and
how they’ve evolved over time. It will also enable us to spot new publishing of denialist claims in
real-time.
@@
Prompt 1: System prompt used for zero-shot classification of claims on the CARDS test dataset. The prompt is framed based on the coding manual retrieved from the supplimantary material of Coan et al. [2021].
@@ A.5 User Prompt
Question: To what claim does the following text
belongs to?
@@
Question: To what claim does the following text
belongs to?
@@ A.6.1 System prompt
misleading claims about climate change in news
media. You are asked to classify whether a text
excerpt belongs to one of the following labels
separated by a comma: 0_0, 1_1, 1_2, 1_3, 1_4,
@@ A.8 Prompt Template for Open Source Models
misleading claims about climate change in news
media. Classify the following text into one of the
27 classes.
Question: To which claim does the following text
belongs to?
@@ A.11 Sample of Classified Claims from Social Media Data
The table below shows a sample of claims sourced from social media data, as described in section 4.3. These claims have been annotated by two climate communication experts using the CARDS taxonomy of false or misleading claims about climate change, as proposed by Coan et al. [2021] (illustrated in Appendix A.1). In addition, the table shows instances of the alignment ([Uncaptioned image]) and misalignment ([Uncaptioned image]) of the GPT-3.5-turbo model – fine-tuned on the expert-annotated dataset from Coan et al. [2021] – and the CARDS model with the expert annotations.

### v2

@@ 4.4 Expert annotation of claims in social media data
Two climate change communication experts, each with over 20 years of experience as professors and academic researchers on how climate science is communicated, annotated the sampled paragraphs from mostly engaging articles on social media (described in section 4.3) on the super-claim and sub-claim levels per the CARDS coding manual[15].
@@ 5.1 Evaluating the performance of LLMs on CARDS data
Using the zero-shot approach (described in section 4.1) to classify climate misinformation on CARDS test dataset, we encountered some inconsistencies in the way some LLMs are responding to our task. Figure 2 in the appendix shows the rate of valid responses per model. The prominent source of this inconsistency is the failure of models to comply with the instructions outlined in the prompt, resulting in responses that are irrelevant or with unstructured formatting. A similar inconsistent behavior of LLMs was also reported in prior research evaluating open-source LLMs on 20 NLP tasks [45]. However, we observed instances of Llama3.1-8B failing to classify claims because it is "not sure if this is the right place to post this [request]" and that the model is "trying to get a better".
@@
Though GPT-4o is a much larger model compared to RoBERTalarge, in terms of parameter size, it is more budget friendly for stakeholders in climate change discourse such as researchers, policymakers, or think tanks to utilize this model when classifying climate misinformation as it does not require substantial investment in computational resources or expertise to develop a customized model such as CARDS for classifying climate misinformation. However, detecting false or misleading claims does not include the models’ capability to respond to or mitigate such claims unless they are augmented with relevant external knowledge as seen in use cases for fact-checking climate change claims[17]. Accordingly, future research is needed to explore and extend our work to determine whether models capable of detecting climate misinformation can augment general models, such as GPT-4o, with information about which claims are false or misleading to help orient these general models toward guard-railing against such claims.
@@
Fine-tuned GPT-3.5-turbo RoBERTalarge
Claim Code Precision Recall F1 Support Precision Recall F1 Support
0 0.94 0.94 0.94 491 0.76 0.96 0.85 491
1 0.71 1.00 0.83 24 0.71 0.62 0.67 24
2 0.93 0.93 0.93 14 1.00 0.79 0.88 14
3 0.57 1.00 0.73 8 0.57 0.50 0.53 8
4 0.95 0.94 0.95 274 0.93 0.63 0.75 274
5 0.99 0.89 0.94 103 0.90 0.62 0.74 103
Accuracy 0.94 914 0.81 914
Macro avg 0.85 0.95 0.88 914 0.81 0.69 0.74 914
Weighted avg 0.94 0.94 0.94 914 0.83 0.81 0.80 914
@@
On the other hand, we observed a poor performance by the fine-tuned model in detecting claims related to climate impacts are not bad. Reviewing the annotated super-claims and sub-claims by experts, we found that the fine-tuned model is unable to accurately classify sub-claims within this category about the impacts of climate change on animal and plant species (see sub-claim 3.2 within the taxonomy of claims in Appendix A.2). We found that the fine-tuned GPT-3.5-turbo is biased toward inaccurately classifying claims regarding climate change impacts on animal and plant species as false, possibly due to biases in the CARDS training data that we originally fine-tuned the model on (see Limitations section). This indicates additional fine-tuning is needed to enhance the ability of the model to differentiate between text describing positive versus negative impacts of climate change.
@@ 6 Conclusion
As developers and researchers test the potential of LLMs to persuade and misinform at scale [47, 8], evaluating the potential for LLMs to be part of the solution for governing online dis/misinformation rather than the problem becomes a task of great importance. In this context, the overarching goal of this paper was to demonstrate the crucial role of human-oversight when leveraging LLMs for governance by evaluating the performance of, both open-source and proprietary, LLMs in the application of classifying false or misleading claims about climate change. We also aimed to compare (1) how these models performed against existing tools designed for the same task, and (2) with the assessments made by climate-communication and mis-information experts.
@@
Another limitation is the reliance of our research on a single taxonomy of claims developed by the authors of CARDS [15] that was leveraged to fine-tune GPT-3.5.-turbo. This introduces two sets of biases. First, as the CARDS annotated dataset was based on an expert review of climate skeptic and contrarian domains, whether the GPT-3.5-turbo model is capable of precisely discriminating between accurate and inaccurate climate change claims within high credible sources (e.g. The New York Times, CNN, etc.) is an open question. One way to address this problem, as the authors of the CARDS model have recently moved toward, is a two-stage approach for classifying claims that first determines the veracity of the claim and in the second stage labels the category of false claims [48]. A next step, therefore, is to benchmark GPT-3.5-turbo model’s performance in classifying claims from high credible sources for comparison to the updated CARDS model and continue fine-tuning as necessary on text sourced from domains with varying credibility.
A second model bias is the inability to classify claims in social media posts that do not fall within the CARDS taxonomy. For instance, the original taxonomy upon which CARDS is based includes claims about human health impacts as a category but was excluded in the final model due it its low prevalence [15]. However, this may be a function of the ideological-skew of climate skeptic blogs on which CARDS was trained that ignored this dimension of climate change impacts and/or temporal trends increasing the prominence of health impacts. As a result, climate change communication experts annotating the social media claims observed a substantial number of false claims about the health impacts of climate change on humans that did not fall within the current CARDS model and which the GPT 3.5-Turbo was unable to classify.
In addition, the model’s poor performance in classifying claims about the impacts of climate change on animal and plant species (see Table 2), could also be attributed to the under-representation of these examples in the CARDS dataset used for fine-tuning. Moving forward, fine-tuning the GPT-3.5-turbo model on additional expert annotated datasets, for example from Climate Feedback  [^10] https://science.feedback.org would likely enhance the model’s performance in accurately classifying a wider range of claims. These limitations stress the importance of benchmarking the performance of LLMs against data collected from “the wild” as we did in this paper and fine-tuning accordingly to ensure optimal performance in detecting misinformation online.
@@ A.1 Climate Change Keywords
A full list of the compiled climate change keywords identified by climate experts that are used to scrape and filter relevant climate change articles: climate change - climate crisis - climate effects - climate hoax - climate policy - climate resilience - climate science - climate summit - global warming - greenhouse gas - greenhouse gases - IPCC - green energy - climate hypocrisy - paris agreement - paris climate - net zero - net-zero - COP26 - climate conversation - climate test - climate gap - climate activists - climate activist - clean energy - climate negotiations - climate deal - green new deal - climate conference - green technology - green tech - climate fearmongering - climate fears - climate anxiety - carbon capture.
@@ A.4 System Prompt
CARDS: Computer Assisted Recognition of Denial and Skepticism, is a machine learning project. Our aim is to train a computer to automatically detect and categorize misinformation about climate change. The end goal is that a computer can look at some text and successfully identify any climate misinformation - and even identify specific denialist claims. If successful, this will enable us to travel back in time and build a history of climate misinformation, including when myths originated and how they’ve evolved over time. It will also enable us to spot new publishing of denialist claims in real-time.
@@
Prompt 1: System prompt used for zero-shot classification of claims on the CARDS test dataset. The prompt is framed based on the coding manual retrieved from the supplimantary material of [15].
@@ A.5 User Prompt
Question: To what claim does the following text belongs to?
@@
Question: To what claim does the following text belongs to?
@@ A.6.1 System prompt
misleading claims about climate change in news media. You are asked to classify whether a text excerpt belongs to one of the following labels separated by a comma: 0_0, 1_1, 1_2, 1_3, 1_4,
@@ A.8 Prompt Template for Open Source Models
misleading claims about climate change in news media. Classify the following text into one of the 27 classes.
Question: To which claim does the following text belongs to?
@@ A.11 Sample of Classified Claims from Social Media Data
The table below shows a sample of claims sourced from social media data, as described in section 4.3. These claims have been annotated by two climate communication experts using the CARDS taxonomy of false or misleading claims about climate change, as proposed by [15] (illustrated in Appendix A.1). In addition, the table shows instances of the alignment ([Uncaptioned image]) and misalignment ([Uncaptioned image]) of the GPT-3.5-turbo model – fine-tuned on the expert-annotated dataset from [15] – and the CARDS model with the expert annotations.